OpenStudy (yamyam70):

Basic trigonometry, I have a solution, want someone to check, I'll type in comment section :)

OpenStudy (yamyam70):



OpenStudy (yamyam70):

The calculator is showing me, 2 - square root of 3 . How do I get there?

OpenStudy (shamim):

multily numerator n denominator by √3-1

OpenStudy (yamyam70):

ohh we need to rationalize, can someone help me with that? :)

OpenStudy (solomonzelman):

so so far you get, -1+√3 ------ 1+√3 re-phrasing, √3 - 1 ----- √3 + 1 then multiply both top and bottom times √3 - 1

OpenStudy (yamyam70):

hmm wait up Ill try it on paper

OpenStudy (solomonzelman):

Sure, go ahead...

OpenStudy (yamyam70):

(square root of 3 -1)^2 is FOIL method right? same as polynomials?

OpenStudy (solomonzelman):

yes. (a-b)² = (a-b)(a-b) = a²-2ab+b² Also, To get √ , Click and hold ALT, click 2 5 1 on the number pad on the right, release ALT. To get ² , Click and hold ALT, click 2 5 3 on the number pad on the right, release ALT. To get ³ , Click and hold ALT, click 0 1 7 9 on the number pad on the right, release ALT.

OpenStudy (yamyam70):

pardon, the typings I use only laptop :) and thank you :)